OpenStudy (deadshot):

What is the 19th term of a geometric sequence where a_1=14 and a_9=358.80 ? Round the common ratio and 19th term to the nearest hundredth.

OpenStudy (campbell_st):

well a term in a geometric series is \[a_{n} = a_{1} \times r^{n - 1}\] so you know the 9th term, use it to find the common ratio \[358.8 = 14 \times r^{9 -1}\] which becomes \[r^8 = \frac{358.8}{14}\] solve for r then you can use the formula above to find the 19th term

OpenStudy (deadshot):

so then r is approximately 1.5?

OpenStudy (campbell_st):

thats correct

OpenStudy (deadshot):

Thank you!